# Weyl law for 1-cycles

### v1

@@ 3.4. Approximating discrete families of cycles
and it might be the case that $\tau(v_{C},v_{E})+\tau(v_{E},w_{E}^{D}(x))\neq\tau(v_{C},w^{E}_{D}(x))$ (if the $\tau_{vw}$ could be chosen so that they verified $\tau(v_{1},v_{2})+\tau(v_{2},v_{3})=\tau(v_{1},v_{3})$ for every triplet of adjacent vertices $v_{1},v_{2},v_{3}\in X_{0}$, then we could construct a filling $G:X_{0}\to\mathcal{Z}_{k}(M)$ of $F$ such that $G(v)-G(w)=\tau(v,w)$ and hence $\operatorname{\mathbf{M}}(G(v)-G(w))\leq\varepsilon$ for any adjacent vertices $v$ and $w$, and that is not possible for an arbitrary $\varepsilon$-fine family $F$ as the one we have). Several attempts to use this approach or a similar one have been made by the author (for example, allowing finite sums of the $\tau_{vw}$) but all of them have failed. Nevertheless, (5) seems to be much more versatile to work with because we do not have the previously mentioned problem with the vertex $v_{C}$ and also it allows to circumvent other technical issues that arises when interpolating in cells of dimension greater than $1$. Let us explain first how that interpolation works.
Given a cell $C$ and $D\in\mathcal{D}_{C}$, let $w^{D}_{C}$ be the vertex of $C$ that minimizes $\operatorname{\mathbf{M}}(F(v)\llcorner D)$ among all $v\in V(C)$. As we want to decrease mass, the strategy is to start from $x\in\partial C$ and replace $F(w^{D}_{C}(x))\llcorner D$ by $F(w^{D}_{C})\llcorner D$ one domain $D\in\mathcal{D}_{C}$ at a time as we approach the center of the cell $C$. This is done by adding at each step $\partial[\tau(w_{C}^{D}(x),w_{C}^{D})]$ to our $k$-cycle, which causes a change in $I_{j}(x)$ by adding $\tau(w_{C}^{D}(x),w_{C}^{D})\llcorner\partial D$ (which is small in mass provided the grid $\mathcal{D}_{C}$ is chosen accurately). When we get to the centre of $C$ we have $w_{C}^{D}(x)=w_{C}^{D}$ for all $D$ by construction. Nevertheless, it might happen that $I_{j}(x_{1})\neq I_{j}(x_{2})$ for some $x_{1},x_{2}\in\partial\operatorname{Center}(C)$. This is due to the fact that in general
@@
Given $\tau_{1},...,\tau_{K}\in\mathcal{I}_{k+1}(M)$ and a collection of balls =$\mathcal{B}=\{B_{1},...,B_{L}\}$ we say that $\mathcal{B}$ is compatible with $\tau_{1},...,\tau_{K}$ if
@@
[…]
@@
$\operatorname{Adj}(v)$ =$=\{w\in X_{0}:\exists C\text{ cell of }X\text{ s.t. }v,w\in C\}$
@@
To each cell of $X$ we associate a grid in the following way. For 00-dimensional cells we just consider $\mathcal{D}_{v}=\{D_{1}^{v},...,D_{L}^{v}\}$ as defined above. If $1\leq\dim(C)\leq p$ we define
@@
    $F_{j}$ is $C(p)c(k,j)(\frac{L}{r})^{n(k,j)}\varepsilon$-fine.
@@
$z_{C}$ $=F(v_{C})+\partial\tau_{C};$
$I_{C}$ $=\sum_{D\in\mathcal{D}_{C}}\tau(v_{C},w^{D}_{C})\llcorner\partial D.$
@@
among all possible values of $F_{j}(x)$ for $x\in C\cap X_{j}(q_{j})_{0}$ according to our inductive hypothesis. Thus when extending $F_{j}$ from $\partial C$ to all of $C$, we will homotop $F_{j}|_{\partial C\cap X_{j}(q_{j})_{0}}$ to $z_{C}$ in a way that decreases (6) as we approach to $\operatorname{Center}(C)$.
@@
$I_{1}(x)=\sum_{D\in\mathcal{D}_{C}}\tau(v_{C},w^{D}_{C}(x))\llcorner\partial D$
@@
$F_{j}(y)-F_{j}(x)$ $=\partial[\tau_{C}(y)-\tau_{C}(x)]$
$=\partial[(\tau(v_{C},w_{C}^{D_{E}}(x))-\tau(v_{C},w_{C}^{D_{E}}(y)))\llcorner D_{E}]$
@@
Inductive step. Assume $F_{j-1}:X_{j-1}(q_{j-1})_{0}\to\mathcal{Z}_{k}(M)$ has been defined satisfying the inductive property. Let $\overline{q}_{j}=3q_{j-1}$. First we will define $\overline{F}_{j}(x)$, $\overline{I}_{j}(x)$ and $\overline{w}^{D}_{C}(x)$ with domain $(C\setminus\operatorname{Center}(C))\cap X(\overline{q}_{j})_{0}$ for each $j$-cell $C$ of $X(q)$. Fix a $j$-cell $C$. For $x\in\partial C\cap X(\overline{q}_{j})_{0}$ we set $\overline{F}_{j}(x)=F_{j-1}(\Xi(x))$, $\overline{I}_{j}(x)=I_{j-1}(\Xi(x))$ and we define $w^{D}_{\partial C}(x)=w^{D^{\prime}}_{E}(x)$ where $E\subseteq\partial C$ is a $(j-1)$-cell of $X$ containing $x$ and $D^{\prime}\in\mathcal{D}_{E}$ is the unique domain such that $D\subseteq D^{\prime}$. The previous is well defined by (7) in the Inductive Property. For $x\in(C\setminus\operatorname{Center}(C))\cap X(\overline{q}_{j})_{0}$ we set
@@
$\overline{F}_{j}(x)=F_{j-1}(\Xi(x))+\partial\big{[}\sum_{\begin{subarray}{c}D\in\mathcal{D}_{C}\\ l(D)\leq d_{\infty}(x,\partial C)\end{subarray}}\tau(w^{D}_{\partial C}(\Xi(x)),w^{D}_{C})\llcorner D\big{]}.$
@@
$\overline{I}_{j}(x)=I_{j-1}(\Xi(x))+\sum_{\begin{subarray}{c}D\in\mathcal{D}_{C}\\ l(D)\leq d_{\infty}(x,\partial C)\end{subarray}}\tau(w^{D}_{\partial C}(\Xi(x)),w^{D}_{C})\llcorner\partial D$
@@
and $\overline{F}_{j}$ is well-defined as it extends $F_{j-1}\circ\Xi:X_{j-1}(\overline{q}_{j})_{0}\to\mathcal{Z}_{k}(M)$ (same for $\overline{I}_{j}$). Consider $E\in\operatorname{Faces}_{j}(X_{j}(\overline{q}_{j}))$ such that $E\subseteq C\setminus\operatorname{Center}(C)$ and let $E^{\prime}\in\operatorname{Faces}_{j-1}(X_{j-1}(q_{j-1}))$ and $C^{\prime}\in\operatorname{Faces}_{j-1}(X)$ be such that $\Xi(E)\subseteq E^{\prime}\subseteq C^{\prime}$ (there could be more than one pair of cells $(E^{\prime},C^{\prime})$ with this property, in that case we just pick one). Denote $l(x)=d_{\infty}(x,\partial C)$, then there exists $l\in\mathbb{N}$ such that if $x\in E$ either $l(x)=l-1$ or $l(x)=l$. Define
$\mathcal{D}_{C}^{E}=\{D\in\mathcal{D}_{C}:D\subseteq D^{\prime}\text{ for some}D^{\prime}\in\mathcal{D}_{C^{\prime}}^{E^{\prime}}\}\cup\{D_{E}\}$
where $D_{E}\in\mathcal{D}_{C}$ is the unique domain which verifies $l(D_{E})=l$. Let $\{B_{i}^{E}\}_{i\in I_{E}}$ be an admissible family which covers $D_{E}\cup\bigcup_{i\in I_{E^{\prime}}}B_{i}^{E^{\prime}}$. Observe that if $x,y\in E$ then
$\mathcal{F}(\overline{F}_{j}(x),\overline{F}_{j}(y))\leq\mathcal{F}(F_{j-1}\circ\Xi(x),F_{j-1}\circ\Xi(y))+\\ \operatorname{\mathbf{M}}\bigg{(}\sum_{\begin{subarray}{c}D\in\mathcal{D}_{C}\\ l(D)\leq d_{\infty}(x,\partial C)\end{subarray}}\tau(w_{\partial C}^{D}(\Xi(x)),w^{D}_{C})\llcorner D-\sum_{\begin{subarray}{c}D\in\mathcal{D}_{C}\\ l(D)\leq d_{\infty}(y,\partial C)\end{subarray}}\tau(w_{\partial C}^{D}(\Xi(y)),w^{D}_{C})\llcorner D\bigg{)}$
The first term is bounded by $C(p)c(k,j-1)(\frac{L}{r})^{n(k,j-1)}\varepsilon$ by inductive hypothesis, and second one is dominated by
$\sum_{D\in\mathcal{D}_{C}}\sum_{v,w\in V(C)}\operatorname{\mathbf{M}}(\tau_{v,w}\llcorner D)=\sum_{v,w\in V(C)}\operatorname{\mathbf{M}}(\tau_{vw})\leq{2^{j}\choose 2}\varepsilon$
hence we have the required $\varepsilon$-fineness for $\overline{F}_{j}$. Therefore properties (B1) to (B6) follow.
@@ 3.5. Approximating discrete families of chains
There exist constants $\varepsilon_{0}(k,p,M)>0$, $\delta_{0}(k,p,M)>0$, $q(k,p,M)\in\mathbb{N}$ and $N(k,p)\in\mathbb{N}$ such that the following is true. Let $G:X_{0}\to\mathcal{I}_{k+1}(M)$ be an $\varepsilon$-fine family, $\varepsilon\leq\varepsilon_{0}$, $10\leq k\leq n-$ 1, $\dim(X)=p$. Let $0<\delta<\delta_{0}$. Then there exists a discrete family $G^{\prime}:X(q)_{0}\to\mathcal{I}_{k+1}(M)$ such that if $F(x)=\partial G(x)$ and $F^{\prime}(x)=\partial G^{\prime}(x)$ then
@@
    $G^{\prime}$ is $c(k,p)(\frac{L}{\delta})^{n(k,p)}\varepsilon$-fine and $(N(k,p),\delta)$-localized.
@@
    $G^{\prime}$ is $c(k,p)(\frac{L}{\delta})^{n(k,p)}\varepsilon$-fine and $(K(N,k,p),K(N,k,p)\delta)$-localized.
@@
and define maps $G_{j}:\tilde{X}_{j}(q_{j})_{0}\to\mathcal{I}_{k+1}(M)$ taking the equivalence relation $\sim$ on $\operatorname{Faces}(\tilde{X})$ into account so that $\varepsilon$-fineness and $\delta$-localization also hold for adjacent vertices in $X(2q_{j}+1)$ which are not neighbours in $\tilde{X}(q_{j})$. The main difference with Proposition 3.17 is that the chains $G(x)$ may have very large mass and then we have to be careful in our construction in order to avoid amplifying the volumes of our chains by big factors. We will do that by using the interpolation techniques from Proposition 3.30.
We define $G_{0}:\tilde{X}_{0}\to\mathcal{I}_{k+1}(M)$ in the a very similar way as in Proposition 3.17. For $v\in X_{0}\subseteq X(3)_{0}=\tilde{X}_{0}$ we set $G_{0}(v)=G(v)$. Given two adjacent vertices $x,y$ in $\tilde{X}_{0}\setminus X_{0}$, let $E$ be the unique cell of $X$ such that $x,y\in\operatorname{Center}(E)$ and denote $v=\Xi(x)$, $w=\Xi(y)$ being $v,w\in V(E)$. As $\mathcal{F}(G(v),G(w))\leq\varepsilon$ and $G$ is $\delta$-localized, there exist $\tau_{vw}\in\mathcal{I}_{k+1}(M)$ and $\sigma_{vw}\in\mathcal{I}_{k+2}(M)$ such that $G(v)-G(w)=\tau_{vw}+\partial\sigma_{vw}$, $\operatorname{\mathbf{M}}(\tau_{vw})+\operatorname{\mathbf{M}}(\sigma_{vw})\leq\varepsilon$ and $\operatorname{spt}(\tau_{vw}),\operatorname{spt}(\sigma_{vw})\subseteq\bigcup_{i\in I_{E}}U_{i}^{E}$. Thus given a face $E$ of $X$ and $x\in V(\operatorname{Center}(E))$ we define
@@
Assume the Inductive Property holds for $j-1$, we sketch how to do the inductive step (obtaining the base case $j=1$ from $G_{0}$ is analog). Let $\mathcal{E}$ be an equivalence class of $j$-faces under $\sim$ and let $E\in\mathcal{E}$. As $R^{q_{j}}F$ is constant on $E$ (because so it is on the top dimensional cells $C_{v}$ of $\tilde{X}$), we know that $\partial G_{j-1}(x)=F(v_{E})$ for every $x\in\partial E\cap\tilde{X}(q_{j-1})_{0}$. Using this, we can interpolate between $\{G(v):v\in V(E)\}$ in the same way as we did between the cycles $\{F(v):v\in V(E)\}$ in the proof of Proposition 3.30, just replacing a “family of cycles” by a “family of chains with constant boundary”. Indeed, the only thing we need to use for such interpolation is that for $x,y$ vertices of some $C_{v}$, the difference $G(x)-G(y)$ is a $(k+1)$-cycle of small flat norm, hence there exists $\sigma_{xy}\in\mathcal{I}_{k+2}(M)$ such that $\partial\sigma_{xy}=G(x)-G(y)$ and $\operatorname{\mathbf{M}}(\sigma_{xy})\leq\varepsilon$. We define the corresponding grids $\mathcal{D}_{E}$ for each cell $E$ of $\tilde{X}$ in the same way as done in the proof of Proposition 3.30 but also requiring that $\mathcal{D}_{E}=\mathcal{D}_{E^{\prime}}$ if $E\sim E^{\prime}$ which can be done in the following way. Consider the equivalence relation $\approx$ in $\tilde{X}_{0}$ given by $v\approx v^{\prime}$ if and only if there exist $E,E^{\prime}\in\operatorname{Faces}(\tilde{X})$ such that $v\in E$, $v^{\prime}\in E^{\prime}$, $E\sim E^{\prime}$ and $T_{EE^{\prime}}(v)=v^{\prime}$. Given an equivalence class $\mathcal{V}$ of $\sim$, we fix a grid $\mathcal{D}_{\mathcal{V}}$ compatible with $\{\sigma(w,w^{\prime}):w,w^{\prime}\in\bigcup_{v\in\mathcal{V}}\operatorname{Adj}(v)\}$ where $\operatorname{Adj}(v)$ is taken in the simplicial complex $\tilde{X}$ and we set $\mathcal{D}_{v}=\mathcal{D}_{\mathcal{V}}$ if $v\in\mathcal{V}$. Then by construction, $\mathcal{D}_{v}=\mathcal{D}_{v^{\prime}}$ if $v\approx v^{\prime}$ and this implies that $\mathcal{D}_{E}=\mathcal{D}_{E^{\prime}}$ if $E\sim E^{\prime}$ following Definition 3.41 (where grids for higher dimensional cells are constructed as the intersection of the grids of their vertices). We will use the operation of adding
@@
to change $w^{D}_{E}(x)=w$ by $w^{D}_{E}(x)=v$ and $I_{j}(x)$ by $I_{j}(x)+\sigma_{vw}\llcorner\partial D$. That way we can extend $G_{j}$, $I_{j}$ and $w^{D}_{E}(x)$ to all of $E$ verifying properties (C1) to (C9) for $j$-faces $F\subseteq E$ following the proof of Proposition 3.30. Given $E^{\prime}\in\mathcal{E}$, $E^{\prime}\neq E$, we extend $\Delta^{j}_{EE^{\prime}}:E\cap\tilde{X}(q_{j})_{0}$ to be constantly equal to $\Delta^{j}_{EE^{\prime}}(v_{E})=G_{0}(v_{E})-G_{0}(T_{EE^{\prime}}(v_{E}))$ on $E\setminus\partial E$. Then by inductive hypothesis, $\operatorname{\mathbf{M}}(\Delta_{EE^{\prime}}^{j}(x))\leq C(j-1)\varepsilon$ for every $x$. We set

### v2

@@ 3.4. Approximating discrete families of cycles
and it might be the case that $\tau(v_{C},v_{E})+\tau(v_{E},w_{E}^{D}(x))\neq\tau(v_{C},w^{E}_{D}(x))$ (if the $\tau_{vw}$ could be chosen so that they verified $\tau(v_{1},v_{2})+\tau(v_{2},v_{3})=\tau(v_{1},v_{3})$ for every triplet of adjacent vertices $v_{1},v_{2},v_{3}\in X_{0}$, then we could construct a filling $G:X_{0}\to\mathcal{I}_{k+1}(M)$ of $F$ such that $G(v)-G(w)=\tau(v,w)$ and hence $\operatorname{\mathbf{M}}(G(v)-G(w))\leq\varepsilon$ for any adjacent vertices $v$ and $w$, and that is not possible for an arbitrary $\varepsilon$-fine family $F$ as the one we have). Several attempts to use this approach or a similar one have been made by the author (for example, allowing finite sums of the $\tau_{vw}$) but all of them have failed. Nevertheless, (5) seems to be much more versatile to work with because we do not have the previously mentioned problem with the vertex $v_{C}$ and also it allows to circumvent other technical issues that arise when interpolating in cells of dimension greater than $1$. Let us explain first how that interpolation works.
Given a cell $C$ and $D\in\mathcal{D}_{C}$, let $w^{D}_{C}$ be the vertex of $C$ that minimizes $\operatorname{\mathbf{M}}(F(v)\llcorner D)$ among all $v\in V(C)$. As we want to decrease mass, the strategy is to start from $x\in\partial C$ and replace $F(w^{D}_{C}(x))\llcorner D$ by $F(w^{D}_{C})\llcorner D$ one domain $D\in\mathcal{D}_{C}$ at a time as we approach the center of the cell $C$. This is done by subtracting at each step $\partial[\tau(w_{C}^{D}(x),w_{C}^{D})]$ to our $k$-cycle, which causes a change in $I_{j}(x)$ by subtracting $\tau(w_{C}^{D}(x),w_{C}^{D})\llcorner\partial D$ (which is small in mass provided the grid $\mathcal{D}_{C}$ is chosen accurately). When we get to the centre of $C$ we have $w_{C}^{D}(x)=w_{C}^{D}$ for all $D$ by construction. Nevertheless, it might happen that $I_{j}(x_{1})\neq I_{j}(x_{2})$ for some $x_{1},x_{2}\in\partial\operatorname{Center}(C)$. This is due to the fact that in general
@@
Given $\tau_{1},...,\tau_{K}\in\mathcal{I}_{k+1}(M)$ and a collection of balls $\mathcal{B}=\{B_{1},...,B_{L}\}$ we say that $\mathcal{B}$ is compatible with $\tau_{1},...,\tau_{K}$ if
@@
$\operatorname{\mathbf{M}}(\tau_{i})\geq\int_{r}^{2r}\operatorname{\mathbf{M}}(\tau_{i}\llcorner\partial B_{s})ds>\int_{A_{i}}K\frac{\operatorname{\mathbf{M}}(\tau_{i})}{r}ds=K\frac{\operatorname{\mathbf{M}}(\tau_{i})}{r}\mathcal{L}^{1}(A_{i})$
@@
$\operatorname{Adj}(v)$ $=\{w\in X_{0}:\exists C\text{ cell of }X\text{ s.t. }v,w\in C\}$
@@
To each cell of $X$ we associate a grid in the following way. For 0-dimensional cells we just consider $\mathcal{D}_{v}=\{D_{1}^{v},...,D_{L}^{v}\}$ as defined above. If $1\leq\dim(C)\leq p$ we define
@@
    $F_{j}$ is $C(p)c(k,j)(\frac{L}{r})^{n(k,j)}\varepsilon$-fine in $X_{j}$.
@@
$z_{C}$ $=F(v_{C})-\partial\tau_{C};$
$I_{C}$ $=-\sum_{D\in\mathcal{D}_{C}}\tau(v_{C},w^{D}_{C})\llcorner\partial D.$
@@
among all possible values of $F_{j}(x)$ for $x\in C\cap X_{j}(q_{j})_{0}$ according to our inductive hypothesis. Thus when extending $F_{j}$ from $\partial C$ to all of $C$, we will homotop $F_{j}|_{\partial C\cap X_{j}(q_{j})_{0}}$ to $z_{C}$ in a way that decreases (6) as we approach $\operatorname{Center}(C)$.
@@
$I_{1}(x)=-\sum_{D\in\mathcal{D}_{C}}\tau(v_{C},w^{D}_{C}(x))\llcorner\partial D$
@@
$F_{j}(y)-F_{j}(x)$ $=\partial[\tau_{C}(x)-\tau_{C}(y)]$
$=\partial[(\tau(v_{C},w_{C}^{D_{E}}(y))-\tau(v_{C},w_{C}^{D_{E}}(x)))\llcorner D_{E}]$
@@
Inductive step. Assume $F_{j-1}:X_{j-1}(q_{j-1})_{0}\to\mathcal{Z}_{k}(M)$ has been defined satisfying the inductive property. Let $\overline{q}_{j}=3q_{j-1}$. First we will define $\overline{F}_{j}(x)$, $\overline{I}_{j}(x)$ and $\overline{w}^{D}_{C}(x)$ with domain $(C\setminus\operatorname{Center}(C))\cap X(\overline{q}_{j})_{0}$ for each $j$-cell $C$ of $X(q)$. Fix a $j$-cell $C$. For $x\in\partial C\cap X(\overline{q}_{j})_{0}$ we set $\overline{F}_{j}(x)=F_{j-1}(\Xi(x))$, $\overline{I}_{j}(x)=I_{j-1}(\Xi(x))$ and we define $w^{D}_{\partial C}(x)=w^{D^{\prime}}_{E}(x)$ where $E\subseteq\partial C$ is a (j1)(j-1)-cell of $X$ containing $x$ and $D^{\prime}\in\mathcal{D}_{E}$ is the unique domain such that $D\subseteq D^{\prime}$. The previous is well defined by (B7) in the Inductive Property. For $x\in(C\setminus\operatorname{Center}(C))\cap X(\overline{q}_{j})_{0}$ we set
@@
$\overline{F}_{j}(x)=F_{j-1}(\Xi(x))-\partial\big[\sum_{\begin{subarray}{c}D\in\mathcal{D}_{C}\\ l(D)\leq d_{\infty}(x,\partial C)\end{subarray}}\tau(w^{D}_{\partial C}(\Xi(x)),w^{D}_{C})\llcorner D\big].$
@@
$\overline{I}_{j}(x)=I_{j-1}(\Xi(x))-\sum_{\begin{subarray}{c}D\in\mathcal{D}_{C}\\ l(D)\leq d_{\infty}(x,\partial C)\end{subarray}}\tau(w^{D}_{\partial C}(\Xi(x)),w^{D}_{C})\llcorner\partial D$
@@
and $\overline{F}_{j}$ is well-defined as it extends $F_{j-1}\circ\Xi:X_{j-1}(\overline{q}_{j})_{0}\to\mathcal{Z}_{k}(M)$ (same for $\overline{I}_{j}$). Consider $E\in\operatorname{Faces}_{j}(X_{j}(\overline{q}_{j}))$ such that $E\subseteq C\setminus\operatorname{Center}(C)$ and let $E^{\prime}\in\operatorname{Faces}_{j-1}(X_{j-1}(q_{j-1}))$ and $C^{\prime}\in\operatorname{Faces}_{j-1}(X)$ be such that $\Xi(E)\subseteq E^{\prime}\subseteq C^{\prime}$ (there could be more than one pair of cells $(E^{\prime},C^{\prime})$ with this property, in that case we just pick one). Denote $l(x)=d_{\infty}(x,\partial C)$, then there exists $l\in\mathbb{N}$ such that if $x\in E$ either l(x)=l1l(x)=l-1 or $l(x)=l$. Define
[…]
where $D_{E}\in\mathcal{D}_{C}$ is the unique domain which verifies $l(D_{E})=l$. Let $\{B_{i}^{E}\}_{i\in I_{E}}$ be an admissible family which covers $D_{E}\cup\bigcup_{i\in I_{E^{\prime}}}B_{i}^{E^{\prime}}$. Notice that under this choice, (B4) and (B5) follow. We now prove the $\varepsilon$-fineness of $\overline{F}_{j}$. Let $C^{\prime}$ be a face of $\partial C$ which contains $\Xi(x)$. Then
$\mathcal{F}(\overline{F}_{j}(x),F(v_{C}))\leq\mathcal{F}(\overline{F}_{j}(x),F_{j-1}(\Xi(x)))+\mathcal{F}(F_{j-1}(\Xi(x)),F(v_{C^{\prime}}))+\mathcal{F}(F(v_{C^{\prime}}),F(v_{C})).$
The last two terms of the RHS are bounded by $C(p)c(k,j-1)(\frac{L}{r})^{n(k,j-1)}\varepsilon$ and $\varepsilon$, and the first term is bounded by
$\sum_{D\in\mathcal{D}_{C}}\operatorname{\mathbf{M}}\big(\tau(w^{D}_{\partial C}(\Xi(x)),w^{D}_{C})\llcorner D\big)$ $\leq\sum_{D\in\mathcal{D}_{C}}\sum_{v,w\in V(C)}\operatorname{\mathbf{M}}(\tau_{v,w}\llcorner D)$
$=\sum_{v,w\in V(C)}\operatorname{\mathbf{M}}(\tau_{vw})$
$\leq{2^{j}\choose 2}\varepsilon.$
Therefore, properties (B1) to (B7) follow.
@@ 3.5. Approximating discrete families of chains
There exist constants $\varepsilon_{0}(k,p,M)>0$, $\delta_{0}(k,p,M)>0$, $C(k,p,M)\in\mathbb{N}$, $n(k,p)\in\mathbb{N}$ and $N(k,p)\in\mathbb{N}$ such that the following is true. Let $G:X_{0}\to\mathcal{I}_{k+1}(M)$ be an $\varepsilon$-fine family, $\varepsilon\leq\varepsilon_{0}$, $0\leq k\leq n-1$, $\dim(X)=p$ and denote $F=\partial G$. Let $0<\delta<\delta_{0}$. Then there exists a discrete family $G^{\prime}:X(q)_{0}\to\mathcal{I}_{k+1}(M)$ for $q=q(k,p,M,\delta)$ such that if we denote $F^{\prime}=\partial G^{\prime}$ then
@@
    $G^{\prime}$ is $c(k,p)(\frac{L}{\delta})^{n(k,p)}\varepsilon$-fine in $X$ and $(N(k,p),\delta)$-localized in $X(q)$.
@@
    $G^{\prime}$ is $c(k,p)(\frac{L}{\delta})^{n(k,p)}\varepsilon$-fine in $X$ and $(K(N,k,p),K(N,k,p)\delta)$-localized in $X(q)$.
@@
[…]
We define $G_{0}:\tilde{X}_{0}\to\mathcal{I}_{k+1}(M)$ in the a very similar way as in Proposition 3.14. For $v\in X_{0}\subseteq X(3)_{0}=\tilde{X}_{0}$ we set $G_{0}(v)=G(v)$. Given two adjacent vertices $x,y$ in $\tilde{X}_{0}\setminus X_{0}$, let $E$ be the unique cell of $X$ such that $x,y\in\operatorname{Center}(E)$ and denote $v=\Xi(x)$, $w=\Xi(y)$ being $v,w\in V(E)$. As $\mathcal{F}(F(v),F(w))\leq\mathcal{F}(G(v),G(w))\leq\varepsilon$ and $F$ is $(N,\delta)$-localized, there exist $\tau(v,w)\in\mathcal{I}_{k+1}(M)$ supported in $\bigcup_{i\in I_{E}}U_{i}^{E}$ such that $\partial\tau(v,w)=F(v)-F(w)$ and $\operatorname{\mathbf{M}}(\tau(v,w))\leq\varepsilon$. Thus given a face $E$ of $X$ and $x\in V(\operatorname{Center}(E))$ we define
@@
[…]
@@
to change $w^{D}_{E}(x)=w$ by $w^{D}_{E}(x)=v$ and $I_{j}(x)$ by $I_{j}(x)+\sigma_{vw}\llcorner\partial D$. That way we can extend $G_{j}$, $I_{j}$ and $w^{D}_{E}(x)$ to all of $E$ verifying properties (C1) to (C9) for $j$-faces $F\subseteq E$ following the proof of Proposition 3.27. Given $^{\prime}\in\mathcal{E}, EE$$^{\prime}\neq E$, we extend $\Delta^{j}_{EE^{\prime}}:E\cap\tilde{X}(q_{j})_{0}$ to be constantly equal to $\Delta^{j}_{EE^{\prime}}(v_{E})=G_{0}(v_{E})-G_{0}(T_{EE^{\prime}}(v_{E}))$ on $E\setminus\partial E$. Then by inductive hypothesis, $\operatorname{\mathbf{M}}(\Delta_{EE^{\prime}}^{j}(x))\leq C(j-1)\varepsilon$ for every $x$. We set